OpenStudy (anonymous):

which of the following is the equation of a line that passes through (-2,-1) and (-4,-3)? y=1/2x + 1, y=x + 1, y =1/2x - 1, or y = x - 1

OpenStudy (anonymous):

Hello, lets first look at what the slop must be. We know to have a slope, we must have at least 2 points, and they must lie on a straight line. Lets call these two points P1, and P2.

OpenStudy (anonymous):

The formula for slope is \[m = \frac{y2-y1}{x2-x1}\]

OpenStudy (anonymous):

So in this case, point P1 has coordinates (x1,y1) = (-2,-1) and point P2 has coordinates (x2,y2)=(-4,-3)

OpenStudy (anonymous):

Plugging those into the slope you get (-3-(-1))/(-4-(-2)) = m = 1

OpenStudy (anonymous):

so it is either choice b or choice d

OpenStudy (anonymous):

to find out which, we must use the point-slope form of the line \[m = (y-y1)/(x-x1) or y = m(x-x1) + y1\]

OpenStudy (anonymous):

oooh my!

OpenStudy (anonymous):

yeah that came out like crap

OpenStudy (anonymous):

its y=m(x-x1) + y1

OpenStudy (anonymous):

now we have m, y1, and x1

OpenStudy (anonymous):

what's m?

OpenStudy (anonymous):

the slope

OpenStudy (anonymous):

which is 1 in this case, using the two points we are given

OpenStudy (anonymous):

ugh im just trying to prepare for this Pert test! trying to make sure i understand EVERYTHING

OpenStudy (anonymous):

ahh, answer should be y =x -1

OpenStudy (anonymous):

Show me why

OpenStudy (anonymous):

So we have 2 points, (-2,-1) and (-4,-3), do you know how to find slope given 2 points?

OpenStudy (anonymous):

I think so.. I kind of remember that

OpenStudy (anonymous):

so if you graph the two points, and connect a line its (rise)/(run) so how far do i have to go in the y direction divided by how far i have to go in the x direction between the two points

OpenStudy (anonymous):

in the y direction I go from (-3 to -1) total distance

OpenStudy (anonymous):

and in the x direction i go from -4 to -2

OpenStudy (anonymous):

or sorry, i had that backwards in the y direction we go from -1 to -3 and from -2 to -4 in x

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

lol, ok let me graph something real quick

OpenStudy (anonymous):